There are 428 days in a year on planet x and 107 days in a year on planet W. How many years are ther on planet w that are equal to one year on planet x.
Answer by RAY100(1637)

If we treat this as a proportion problem
.
(428 days / yr ) / 1 yr = ( 107 days / yr ) / Z,,,, where z is corresponding years on W
.
cross multiply and simplify
.
428 *z = 107
.
z = 107 / 428 = 1/4 yr
.
checking
ratio of days per year is 107/428 = 1/4
ratio of years = 1/4 ,,,,,,, ok